OpenStudy (owlet):

Solve all values of x in the interval [0, 2pi] that satisfy the inequality: sin =<1/2

OpenStudy (owlet):

\[\sin \le 1/2\]

OpenStudy (anonymous):

from this graph  sin x = 1/2 x = pi/6 from 0 to pi/6, the graph is less than 1/2, right? so now you have to find the other interval where the graph is below 1/2 :) hint:  hope this helps